OpenStudy (anonymous):

12x^2 + 14x + 12 = 18

OpenStudy (uri):

Arrange the eq first 12x^2 + 14x + 12 = 18 12x^2+14x+12-18 12x^2+14x-6=0

OpenStudy (abb0t):

what are you being asked to do?!

OpenStudy (uri):

Now factorize.

OpenStudy (uri):

12x^2+14x-6 2(6x^2+7x-3)

hero (hero):

Don't forget to set it equal to zero.

hero (hero):

If you do then you can just divide both sides by 2 and you'll have: 6x^2 + 7x - 3 = 0

hero (hero):

And you can finish factoring from there.

OpenStudy (uri):

Now i think we have to use Quadratic formula?